OpenStudy (anonymous):

f(z)=(3*z+1)/(z+4*i) f(2+1)= ? Please help,I am founding fault result.

OpenStudy (anonymous):

f(3)=(3*3+1)/(3+4i)=10(3-4i)/25=6/5+4i/25

OpenStudy (anonymous):

was it f(2+i) or f(2+1) = f(3)?

OpenStudy (anonymous):

Sorry! it's f(2+i)!

myininaya (myininaya):

factorial thats scary

myininaya (myininaya):

i know thats an exclamation mark

myininaya (myininaya):

\[f(2+i)=\frac{3(2+i)+1}{2+i+4i}\] what exactly do you want? you seem to be in upper level class so what should i write this as or i mean what is the goal

OpenStudy (anonymous):

scanning solution now, answer in sec....then breakfast >.>

myininaya (myininaya):

im gonna have a fruit smoothie joe what state do you live in?

OpenStudy (anonymous):

Texas (sadly)

myininaya (myininaya):

i live next door

OpenStudy (anonymous):

solution, might be wrong, i think its right though. hunger has adverse effects on math.

myininaya (myininaya):

so im sadder

OpenStudy (anonymous):

I am founding (7+3i)/(2+5i) but the answer is 1-i.

OpenStudy (anonymous):

next door? Louisiana? i think thats spelled right

myininaya (myininaya):

no arkansas

OpenStudy (anonymous):

LOL i added wrong. 14+15 is 29, not 39

myininaya (myininaya):

see we arent all hicks

OpenStudy (anonymous):

you are on the right path alexis, you just need to get the complex number out of the denominator of that fraction.

OpenStudy (anonymous):

thats true, we all arent hick lolol

OpenStudy (anonymous):

I was never think do mutliply with (2+5i). Thanks a lot guys! :D

myininaya (myininaya):

\[\frac{7+3i}{2+5i}*\frac{2-5i}{2-5i}=\frac{14-35i+6i-15i^2}{4-5i^2}\]

myininaya (myininaya):

lol

myininaya (myininaya):

go eat your breakfast

OpenStudy (anonymous):

alright.....breakfast, i'll see you guys later lol

OpenStudy (anonymous):

I hope you will be here to answer and at my next questions :D lol

myininaya (myininaya):

are you in algebra?

myininaya (myininaya):

im confused because some of these questions seems algebra like and some look little different

OpenStudy (anonymous):

Yes,Algebra. Coplex part.

myininaya (myininaya):

complex part?

OpenStudy (anonymous):

At complex chapter.

OpenStudy (anonymous):

still up

myininaya (myininaya):

yep my hours are messed up

OpenStudy (anonymous):

wow you should get some sleep. not good to stare at computer all night

myininaya (myininaya):

i watched the racoons in backyard at 5 ish

OpenStudy (anonymous):

raccoons in the back yard?!

myininaya (myininaya):

i put cat food out for the wild cats but i seen like 5 racoons gang bang it this morning

myininaya (myininaya):

they are interesting

OpenStudy (anonymous):

Answer to my new question :D

OpenStudy (anonymous):

wild cats! raccoons! what in the world? any opossums?

myininaya (myininaya):

k

OpenStudy (anonymous):

i would think the cats would scare off the raccoons

myininaya (myininaya):

yes i see opossums around here

OpenStudy (anonymous):

i thought you were on the west coast. shows you what i know

OpenStudy (anonymous):

you doing complex variables problems?

OpenStudy (radar):

Alexis please review the multiplication of the denominator: (2+5i)(2-5i). It is not:\[4-5i ^{2}\]

myininaya (myininaya):

because its 4-25i^2 lol